OpenStudy (anonymous):

Find the equilibrium concentrations of A, B, and C for a=1, b=1, and c=2. Assume that the initial concentrations of A and B are each 1.0 M and that no product is present at the beginning of the reaction. Consider the following reaction and associated equilibrium constant: aA(g)+bB(g)⇌cC(g), Kc = 4.0

OpenStudy (anonymous):

So I set it up this way \[4.0 = x ^{2} /(1-x)^{2} \] = 3x^2 -8x+4 =0 I then use the quadratic equation and get x= 2/3 1- (2/3) = .67 but apparently that's not the answer.

OpenStudy (anonymous):

1- (2/3) = .33